OpenStudy (anonymous):

What is the energy of a single photon of infrared light with a wavelength of 1.00 μm? 1 μm = 10^-6 m 1.98 × 10^-19 J 4.29 × 10^-20 J 4.98 × 10^-19 J 5.05 × 10^18 J

OpenStudy (p0sitr0n):

E=hf, h is planck constant E=hc/wavelength use wavelength = 1*10^-6 h = 6.626*10^-34 J/Hz

OpenStudy (anonymous):

that is all

OpenStudy (anonymous):

*is that all?